A 30 cm diameter pipe carries oil of sp. gr. 0.8 at a velocity of 2 m/s. At another section the diameter is 20 cm. Find the velocity at this section and also mass rate of flow of oil.
1
Expert's answer
2020-12-08T07:47:45-0500

Diameter of pipe 1 "d_1=30 cm=0.3\ m"

C/S Area of pipe 1 "A_1 =\dfrac{\pi(d_1)^2}{4}=0.0225\pi\ m^2"


Diameter of pipe 2 "d_2=20\ cm=0.2\ m"

C/S Area of pipe 2 "A_2 =\dfrac{\pi(d_2)^2}{4}=0.01\pi\ m^2"


Velocity at pipe 1 "v_1=2\ m/s"

Velocity at pipe 2 "v_2=?"


Discharge through pipe 1 "Q_1=A_1v_1=0.0225\pi\ m^2 (2\ m/s)"

"=0.045\pi\ m^3/s=0.1414\ m^3/s"


From continuity "Q_1=Q_2+Q_3"

Discharge through pipe 2 "Q_2=A_2v_2=0.01\pi\ m^2 (v_2)"

"=Q_1=0.045\pi\ m^3/s"


Velocity at pipe 2 "v_2=\dfrac{Q_2}{A_2}=\dfrac{0.045\pi\ m^3/s}{0.01\pi\ m^2 }=4.5\ m/s"

"v_2=4.5\ m/s"

Mass rate of flow 


"\dfrac{\Delta m}{\Delta t}=\rho Av=0.8\cdot1000kg/m^3(0.045\pi\ m^3/s)"

"=113.097 kg\/s"

"\dfrac{\Delta m}{\Delta t}=113.097 kg/s"